OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + y = 2 3x + 2y = 9

OpenStudy (anonymous):

from first equation....... y=x-2 put in 2nd equation:; 3x+2(x-2)=9 3x+2x-4=9 5x=13 x=13/5 so,y=(13/5)-2 =3/5

OpenStudy (campbell_st):

slight correction from the 1st equation the equivalent to x is x = 2 - y then substitute into the 2nd equation 3(2 - y) + 2y = 9 6 - 3y + 2y = 9 -y + 6 = 9 solve for y

OpenStudy (usukidoll):